### v1

# A Hierarchical Bin Packing Framework with Dual Manipulators via Heuristic Search and Deep Reinforcement Learning

Abstract
We address the bin packing problem (BPP), which aims to maximize bin utilization when packing a variety of items. The offline problem, where the complete information about the item set and their sizes is known in advance, is proven to be NP-hard. The semi-online and online variants are even more challenging, as full information about incoming items is unavailable. While existing methods have tackled both 2D and 3D BPPs, the 2D BPP remains underexplored in terms of fully maximizing utilization. We propose a hierarchical approach for solving the 2D online and semi-online BPP by combining deep reinforcement learning (RL) with heuristic search. The heuristic search selects which item to pack or unpack, determines the packing order, and chooses the orientation of each item, while the RL agent decides the precise position within the bin. Our method is capable of handling diverse scenarios, including repacking, varying levels of item information, differing numbers of accessible items, and coordination of dual manipulators. Experimental results demonstrate that our approach achieves near-optimal utilization across various practical scenarios, largely due to its repacking capability. In addition, the algorithm is evaluated in a physics-based simulation environment, where execution time is measured to assess its real-world performance.
Note to Practitioners
Bin packing is a key task in logistics and manufacturing, where items must be packed tightly into limited space within a short amount of time. Traditional systems typically use a single robot arm and pack items in the order they arrive, without the ability to rearrange previous placements. This limits space use and makes it hard to improve the result once early placements are made. This work introduces a packing system that incorporates a repacking strategy, allowing previously placed items to be moved when necessary to improve space usage. The system also supports dual-arm operation, where two robot arms work in parallel to reduce overall task time. Items are assigned based on reachability, and placement decisions are made using a combination of predefined rules and AI. The system is intended for environments such as automated warehouses and production lines, where items arrive one by one and must be packed accurately in real time. It was tested in simulation using UR5e robots and open-source planning tools, demonstrating effective use of bin space and faster task execution with the dual-arm setup compared to the single-arm setup.
Index Terms:
Bin packing, dual manipulators, warehouse automation, hierarchical planning
I Introduction
The bin packing problem (BPP) has found applications in delivery, warehouse automation, and manufacturing. BPP is a classic combinatorial optimization problem that aims to pack items into a bin while minimizing wasted space. As BPP is $\mathcal{NP}$-hard [1], various efficient methods have been developed, including heuristics [2, 3, 4, 5, 6]. More recently, reinforcement learning (RL) approaches [7, 8, 9, 10] have demonstrated benefits in reducing computational costs.
Figure 1: Bin packing systems in six different scenarios. (a–c) illustrate configurations with a single manipulator, whereas (d–f) depict those involving dual manipulators. The number of known items (in green and red) varies across scenarios, and the number and placement of manipulators determine which items are accessible.
BPP includes offline, online, and semi-online variants, depending on the availability of item information [11]. While the offline BPP assumes full knowledge of all items, real-world applications typically involve online or semi-online settings, where the information of items are revealed sequentially. Semi-online BPPs are more common in industrial contexts such as logistics and manufacturing. In such settings, strategies for managing uncertainty, such as repacking, can improve adaptability to changes in item arrivals. BPP is also categorized by packing dimension (e.g., 2D, 3D) [12]. Although there has been a line of research on both 2D BPP and 3D BPP, 2D BPP remains underexplored, particularly in achieving near-optimal packing through rotation and/or repacking.
Most existing BPP methods assume a single-robot setting, although coordinating multiple manipulators can significantly reduce execution time through parallelization. Coordinating multiple robots for BPP remains computationally challenging due to the inherent complexity of multi-robot task allocation and multi-arm motion planning [13, 14, 15, 16].
This paper addresses the 2D bin packing problem (BPP) under practical scenarios involving varying numbers of known items, accessible items, and manipulators, as illustrated in Fig. 1. To handle such diverse configurations and support repacking for improved bin utilization, we develop a hierarchical approach that integrates deep reinforcement learning (DRL) with heuristic search. The low-level DRL agent, trained with Asynchronous Advantage Actor-Critic (A3C) [17], selects item positions based on a given bin and item state. On top of this, the high-level search constructs a tree by virtually packing items at positions proposed by the low-level agent, exploring different packing orders and orientations. Unpacking and repacking can be considered to improve bin utilization. A best task sequence is then selected from the candidate branches in the tree and translated into executable actions through integrated task and motion planning, ensuring physical feasibility in both single- and dual-arm settings. In dual-arm scenarios, items are assigned to manipulators based on accessibility, and coordinated actions are planned for parallel execution using off-the-shelf motion planners.
To the best of our knowledge, this is the first work to solve the 2D BPP using a dual-manipulator system under diverse industrial conditions described in Fig. 1. The main contributions of this paper are:
  • A DRL framework that outperforms the state-of-the-art 2D online BPP methods with significantly higher bin utilization.
  • A heuristic search algorithm that supports a wide range of 2D BPP scenarios, including offline, semi-online, and online settings, with varying levels of item accessibility and information, and even multi-manipulator configurations.
  • A repacking strategy that consistently achieves near-100% bin utilization.
II Related Works
Prior work has addressed offline and online 2D BPPs, with recent advances extending to semi-online settings and incorporating rotation and repacking techniques. This section reviews related research on these BPP variants and strategies for rotation and repacking. We also discuss approaches for coordinating multi-robot operations in shared workspaces.
In the offline BPP where all items are known, metaheuristic approaches such as genetic algorithms [3], Tabu Search [6], and layer-building strategies [18] have been proposed. A hybrid DRL-heuristic method [8] also aims to minimize item surface area. However, these methods often suffer from limited generalizability [3], sensitivity to initialization and local optima [6], reliance on handcrafted rules [18], or lack of interaction between learning and rule-based components [8].
In online BPP where items arrive sequentially, heuristic methods such as Skyline [4] and ShelfNextFit [5] have been proposed. Some account for item orientations [9] but generalize poorly across varying item sizes. DRL-based methods improve generalization [7, 19, 10] but cannot revise past decisions, limiting near-optimal bin utilization.
The semi-online BPP, where some future items are partially known, has drawn attention as it reflects realistic industrial scenarios. While packing the current accessible item, partial knowledge of upcoming items is available. In [20], a bounded lookahead heuristic estimates processing costs based on known future items. BPP-$k$ [21] extends an RL model trained on BPP-1 to look $k$ steps ahead via tree search. Other works [22, 23] also employ this approach. However, these approaches do not consider robot accessibility, leaving executability unverified. Moreover, none consider repacking placed items.
To achieve high bin utilization, repacking has emerged as a crucial strategy. [24] integrates heuristic rules and DRL: a heuristic determines whether to pack or unpack based on wasted volume, and the RL agent selects the position. [25] compares values from dual critic networks to decide between packing and unpacking. However, both [24, 25] are limited by myopic item selection (e.g., random or first-in-last-out) without global optimization. Additionally, [25] compares unpacked and packed states unfairly, since the unpacked state has fewer packed items and thus an advantage in value.
While multi-robot manipulation improves packing throughput, it introduces coordination challenges. Some works have addressed item rearrangement in shared workspaces [14, 15, 16], but coordinated bin packing remains largely unexplored.
To fill this gap, we propose a framework for 2D BPP under realistic multi-robot settings. Our method considers item accessibility through forward simulation to ensure feasibility, and integrates coordinated task planning and repacking within a unified hierarchical structure. Leveraging a dual-manipulator system, we enable parallel execution while ensuring packing decisions are both feasible and optimized for bin utilization.
III Problem Formulation
(a) The workspace and cache of the dual manipulators
(b) The bin and an item
(c) The representation of the bin and the item
Figure 2: Illustration of the Bin Packing system. (a) The manipulator is denoted by $m$, $\mathcal{Z}$ represents the workspace of the manipulator, and $C$ is the temporary storage area for unpacked items. (b) The bin is an open cuboid, and the next item $o$ is a solid cuboid. (c) The bin is modeled as a padded binary grid matrix $B$, where each cell corresponds to a low-level position action indexed by its coordinate in the image coordinate system. The values increase row-wise from left to right and top to bottom, with the origin located at the top-left corner. The item is encoded as the rotated vector $\mathbf{l}(\phi)$, and a checkerboard-patterned pixel indicates the top-left corner of the placed item.
Our 2D BPP is formulated as a hierarchical Markov Decision Process (MDP). The high-level MDP focuses on finding a sequence of task primitives determining what to pack or unpack, based on the current bin and items. The low-level MDP determines the precise positions to execute these primitives. After each high-level action is executed, the system transitions to the next state with an updated bin configuration.
As shown in Figs. 2(a) and 2(b), we consider a 2D BPP environment where the bin is modeled as an open cuboid and items as closed cuboids, all projected onto the $xy$-plane. We assume up to two robotic manipulators, denoted by $m_{i}$ for $i\in\{1,2\}$, each with a reachable workspace $\mathcal{Z}_{i}\subset\mathcal{W}\subseteq\mathbb{R}^{3}$. Items arrive sequentially via a conveyor belt (or other similar facilities), which advances as items are removed. The bin and each item $o$ are discretized into grid cells: the bin is a $W\times H$ grid, and item $o$ occupies $w_{o}\times h_{o}$ cells, where $W,H,w_{o},h_{o}\in\mathbb{Z}^{+}$. The bin occupancy is represented by a padded binary matrix $B\in\{0,1\}^{(W+2)\times(H+2)}$, where 00 and $1$ indicate free and occupied cells. The outermost rows and columns are padded with $1$s to represent the bin boundaries (see Fig. 2(c)). The rotated size vector $\mathbf{l}_{o}(\phi)=[l_{x}^{(o,\phi)},l_{y}^{(o,\phi)}]^{T}$ defines the dimensions of item $o$ after rotation by $\phi\in\{0,\frac{\pi}{2}\}$. For example, if $w_{o}=3$ and $h_{o}=1$, then $\mathbf{l}_{o}(0)=[3,1]^{T}$ and $\mathbf{l}_{o}(\frac{\pi}{2})=[1,3]^{T}$. We denote $\mathcal{N}$ as items on the conveyor, $\mathcal{C}$ as unpacked items, and $\mathcal{I}$ as packed items. Each manipulator $m_{i}$ has a buffer space $C_{i}\subset\mathcal{Z}_{i}$ to temporarily store items in $\mathcal{C}$.
The High-level MDP
The high-level MDP governs overall task planning, which is defined as $(\mathcal{S}_{\text{high}},\mathcal{A}_{\text{high}},P_{\text{high}},R_{\text{high}})$. Each state $s_{\text{high}}\in\mathcal{S}_{\text{high}}$ comprises the bin occupancy $B$, the set of items to be packed $\mathcal{N}$, the packed items $\mathcal{I}$, and the unpacked items $\mathcal{C}$:
$s_{\text{high}}=\{B,\mathcal{N},\mathcal{I},\mathcal{C}\}.$
An action $a_{\text{high}}\in\mathcal{A}_{\text{high}}$ is a sequence of task primitives: pack, unpack, or terminate. The pack and unpack primitives have parameters which depend on specific items, while terminate ends the current episode when no further action is necessary or feasible.
The primitive $\texttt{pack}(o,\phi_{o},x_{o},y_{o})$ places item $o$ from $\mathcal{N}$ or $\mathcal{C}$ into the bin with its top-left corner at $(x_{o},y_{o})$ and orientation $\phi_{o}$. While $\phi_{o}$ is selected by the high-level MDP, $(x_{o},y_{o})$ is determined by the low-level MDP. This operation updates the bin occupancy by setting the occupied cells to $1$:
$B^{\prime}(x,y)=1$ (1)
for all $x\in[x_{o}+1,x_{o}+l_{x}^{(o,\phi_{o})}]$ and $y\in[y_{o}+1,y_{o}+l_{y}^{(o,\phi_{o})}]$. The item $o$ is then removed from $\mathcal{N}$ or $\mathcal{C}$ and added to $\mathcal{I}$.
The primitive $\texttt{unpack}(o)$ removes item $o$ from the bin, setting its previously occupied cells to:
$B^{\prime}(x,y)=0$
for all $x\in[{px}_{o}+1,{px}_{o}+l_{x}^{(o,{\psi_{o}})}]$ and $y\in[{py}_{o}+1,{py}_{o}+l_{y}^{(o,\psi_{o})}]$, where $({px}_{o},{py}_{o})$ and $\psi_{o}$ denote the position and orientation of $o$ before unpacking. The item is removed from $\mathcal{I}$ and added to $\mathcal{C}$.
A complete action is composed by sequencing these primitives. For example:
$a_{\text{high}}=($ $\texttt{unpack}(o_{1}),\texttt{unpack}(o_{2}),$
$\texttt{pack}(o_{1},\phi_{o_{1}},x_{o_{1}},y_{o_{1}}),\texttt{pack}(o_{2},\phi_{o_{2}},x_{o_{2}},y_{o_{2}}),$
[…]
unpacks $o_{1}$ and $o_{2}$, then repacks them at new positions, followed by termination. An action may consist solely of terminate, or include only pack.
The high-level reward $r_{\text{high}}$ is defined as the total change in bin occupancy resulting from the executed primitives in $a_{\text{high}}$:
$r_{\text{high}}(s_{\text{high}},a_{\text{high}})=\sum_{\lambda\in a_{\text{high}}}c_{o}w_{o}h_{o},$
where $c_{o}=1$ if $\lambda$ is pack, $c_{o}=-1$ if $\lambda$ is unpack, and $c_{o}=0$ otherwise. The cumulative high-level reward indicates the bin utilization ratio, encouraging space-efficient packing in real-world applications.
The Low-level MDP
As mentioned, the primitive $\texttt{pack}(o,\phi_{o},x_{o},y_{o})$ requires a specific position $(x_{o},y_{o})$ to place item $o$ in the bin. When constructing $a_{\text{high}}$, the low-level MDP $(\mathcal{S}_{\text{low}},\mathcal{A}_{\text{low}},P_{\text{low}},R_{\text{low}},\gamma_{\text{low}})$ is used to determine this position. The low-level state for item $o$ under orientation $\phi\in\{0,\frac{\pi}{2}\}$ is defined as:
$s_{\text{low}}^{(o,\phi)}=\{B,\mathbf{l}_{o}(\phi)\},$
where $B$ is the current bin occupancy and $\mathbf{l}_{o}(\phi)$ is the rotated dimension of $o$. Given $s_{\text{low}}^{(o,\phi)}$, the low-level action $a_{\text{low}}^{(o,\phi)}\in\mathcal{A}_{\text{low}}$ specifies the position for placing the top-left corner of $o$ in the bin. The high-level MDP evaluates both orientations and selects one along with its corresponding position to complete the pack primitive. The bin position is represented by an index from 00 (top-left) to $W\cdot H-1$ (buttom-right), following a row-wise order as illustrated in Fig. 2(c). Given a position $(x_{o},y_{o})$ for item $o$, the low-level action is defined as:
$a_{\text{low}}^{(o,\phi)}=x_{o}+y_{o}\cdot W.$
If no valid position is available, a special no-position action is selected:
$a_{\text{low}}^{(o,\phi)}=W\cdot H.$ (2)
Thus, the low-level action space is defined as $\mathcal{A}_{\text{low}}=\{0,1,\dots,W\cdot H\}$. For each $\phi\in\{0,\frac{\pi}{2}\}$, $a_{\text{low}}^{(o,\phi)}$ is chosen in state $s_{\text{low}}^{(o,\phi)}$ to place item $o$ of dimension $l_{x}^{(o,\phi)}\times l_{y}^{(o,\phi)}$.
The low-level reward is defined as the number of occupied cells adjacent to the packed item $o$:
$r_{\text{low}}(s_{\text{low}}^{(o,\phi)},a_{\text{low}}^{(o,\phi)})=$
$+\sum_{j=1}^{l_{x}^{(o,\phi)}}\left(B(x_{o}+j,y_{o})+B(x_{o}+j,y_{o}+l_{y}^{(o,\phi)})\right)$
$+\sum_{j=1}^{l_{y}^{(o,\phi)}}\left(B(x_{o},y_{o}+j)+B(x_{o}+l_{x}^{(o,\phi)},y_{o}+j)\right).$ (3)
A higher reward is given when $o$ is placed adjacent to more occupied cells; zero reward is given otherwise. The low-level MDP aims to maximize the expected cumulative reward:
$\sum_{j=0}^{\infty}\gamma_{\text{low}}^{j}\,r_{\text{low}}(s_{\text{low}}^{(o_{j},\phi)},a_{\text{low}}^{(o_{j},\phi)}),$
where the discount factor $\gamma_{\text{low}}$ controls the contribution of future rewards. This objective encourages policies that promote compact placements to improve long-term bin utilization.
IV Methods
The overall framework, illustrated in Fig. 3, consists of three main modules: a hierarchical algorithm, task planning, and motion planning.
To solve the 2D BPP, we propose a hierarchical algorithm that integrates DRL with heuristic search (Fig. 5). This hierarchical structure enables interaction between the high- and low-level MDPs through a tree search process. At the low level (Sec. IV-A), an RL agent generates positional actions that determine precise placement positions within the bin. At the high level (Sec. IV-C), the controller operates in two stages. In the first stage, a Depth-First Selective Beam Search (DFS-BS) algorithm expands a search tree by recursively packing items at positions proposed by the low-level agent, while exploring various packing orders and orientations. DFS-BS controls the branching factor by selecting the most promising candidates based on reward values and item sizes. When necessary, the search performs unpacking operations and considers repacking to improve bin utilization. In the second stage, each candidate packing sequence (i.e., each branch of the tree) is evaluated via forward simulation using a reward-based heuristic score, and the highest-scoring sequence is selected as the high-level action. Each high-level action either specifies a set of items to unpack or determines a set of items to pack, including their placement positions and orientations. Executing a high-level action transitions the environment to a new state with updated bin occupancy and item sets.
Subsequently, the task planning module allocates the task primitives from the resulting high-level actions to individual robots and translates them into atomic actions (e.g., pick and place). The motion planning module then generates joint trajectories for each robot to execute the atomic actions in a physically feasible manner.
Figure 3: Overall framework. Our bin packing system integrates a hierarchical algorithm with task and motion planning.
IV-A Low-Level: Position Selection
Figure 4: Actor-critic framework. The input consists of the current bin configuration $B$ and the size vector $\mathbf{l}(\phi)$ of the item.
The low-level MDP is solved by an RL agent that determines the packing position for each item. The agent employs A3C, an actor-critic algorithm, to learn the optimal policy $\pi_{\text{low}}$, which generates $a_{\text{low}}^{(o,\phi)}$ in $s_{\text{low}}^{(o,\phi)}$ for item $o$ with candidate orientation $\phi$, as defined in (5). A3C enables asynchronous training across multiple environments and leverages a value function to evaluate states for improved decision-making. The actor-critic framework is illustrated in Fig. 4.
IV-A1 State input
The agent takes $B$ and $\mathbf{l}_{o}(\phi)$ as input. A convolutional neural network (CNN) encodes $B$ into a feature vector, concatenated with $\mathbf{l}_{o}(\phi)$ and passed through a linear layer. The actor outputs action probabilities $\Pi(a\mid s^{(o,\phi)}_{\text{low}})$, and the critic estimates the state value $V(s^{(o,\phi)}_{\text{low}})$.
IV-A2 Feasibility mask
The placement of item $o$ must consider both the bin boundary and existing packed items. While $o$ should occupy free grid cells within the bin, it is also important to avoid selecting the no-position action when sufficient space exists. To enforce this, we compute a feasibility mask $\mathbf{b}$ at each state before action selection: $\mathbf{b}=[b_{0},b_{1},\ldots,b_{W\cdot H}],$ where $b_{j}=0$ if action $a_{\text{low}}=j$ violates placement constraints, and $b_{j}=1$ otherwise. Inspired by [25], we apply the mask $\mathbf{b}$ to the pre-softmax logits $\mathbf{z}=[z_{0},z_{1},\ldots,z_{W\cdot H}]$ from the linear layer: $\mathbf{z}^{\prime}=\mathbf{b}\odot\mathbf{z}+(1-\mathbf{b})\odot(-10^{8}),$ where $\odot$ denotes element-wise multiplication. This effectively suppresses actions that violate placement constraints, including the no-position action when sufficient space is available, by assigning large negative scores. The final policy is defined via softmax:
$\pi(a_{t}\mid s_{t})=\text{softmax}(\mathbf{z}^{\prime}).$ (4)
IV-A3 Loss function
The A3C algorithm employs parallel actor-learners with a shared global network. Each actor updates the policy $\pi(a_{t}|s_{t};\theta)$ and value function $V(s_{t};\theta_{v})$ using local rollouts. The policy is trained via the policy gradient:
$\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi_{\theta}}\left[\nabla_{\theta}\log\pi(a_{t}|s_{t};\theta)A(s_{t},a_{t})\right],$
where the advantage $A(s_{t},a_{t})=Q(s_{t},a_{t})-V(s_{t})$. The actor and critic losses are defined as:
$L^{\text{actor}}=-\log\pi(a_{t}|s_{t};\theta)A(s_{t},a_{t}),$
$L^{\text{critic}}=\left(R_{t}+\gamma V(s_{t+1})-V(s_{t})\right)^{2}.$
The total loss $L_{\text{total}}$ is the sum of both. Gradients are asynchronously aggregated to update $\theta$ and $\theta_{v}$.
IV-B Interaction Between High-level and Low-level
We introduce a tree $\mathcal{T}=(\mathcal{V},\mathcal{E})$ that unifies the high-level and low-level MDPs, where $\mathcal{V}$ and $\mathcal{E}$ denote the sets of vertices and edges, respectively. The tree is used at the high level to evaluate permutations of task primitives along with their parameters. For each high-level state $s_{\text{high}}=\{B,\mathcal{N},\mathcal{I},\mathcal{C}\}$, a tree $\mathcal{T}$ is constructed, where each vertex $v\in\mathcal{V}$ represents a high-level state of the same form. Using the low-level policy $\pi_{\text{low}}$, a low-level action $a_{\text{low}}^{(o,\phi)}$ is generated for item $o$ with orientation $\phi$, producing a child vertex:
$a_{\text{low}}^{(o,\phi)}=\pi_{\text{low}}(s_{\text{low}}^{(o,\phi)})=\pi_{\text{low}}(\{B,\mathbf{l}_{o}(\phi)\}).$ (5)
This leads to a new vertex $v^{\prime}=\{B^{\prime},\mathcal{N}^{\prime},\mathcal{I}^{\prime},\mathcal{C}^{\prime}\}\in\mathcal{V}$ in the tree, where $B^{\prime}$ is updated via (1), and item $o$ is moved from $\mathcal{N}$ or $\mathcal{C}$ to $\mathcal{I}$ [^1] Item $o$ is removed from $\mathcal{N}$ or $\mathcal{C}$, whichever contains it, and added to $\mathcal{I}$. If $a_{\text{low}}^{(o,\phi)}$ is the no-position action (2), no child is generated and $v$ becomes a leaf vertex. A leaf represents a state where either (i) no items remain in $\mathcal{N}\cup\mathcal{C}$, or (ii) $o$ cannot be placed due to collisions with packed items in $\mathcal{I}$.
From a sequence of vertices from the root of $\mathcal{T}$ and to a leaf, we extract a tuple $(o,\phi_{o},a_{\text{low}}^{(o,\phi)},d(v^{\prime}))$ for each child $v^{\prime}$ of $v$, where $d(v^{\prime})=d(v)+1$ denotes the depth of $v^{\prime}$ in $\mathcal{T}$. Excluding the root, we construct a sequence of such tuples:
$\chi=$ $\big{(}(o_{1},\phi_{o_{1}},a_{\text{low}}^{(o_{1},\phi_{o_{1}})},1),(o_{2},\phi_{o_{2}},a_{\text{low}}^{(o_{2},\phi_{o_{2}})},2),$ (6)
$\dots,(o_{n},\phi_{o_{n}},a_{\text{low}}^{(o_{n},\phi_{o_{n}})},n)\big{)},$
which serves as the basis for generating a high-level action in the form of a task primitive sequence.
[…]
IV-C High-Level: Tree Expansion and Action Generation
Figure 5: Hierarchical algorithm. Each candidate sequence $\chi^{\prime}$ is reordered into $\tilde{\chi}$, evaluated via forward simulation, and the best one is selected to generate the high-level action $a_{\text{high}}$.
We present the high-level procedure of the hierarchical bin packing algorithm, from tree construction to the generation of the final task sequence. The overall process is illustrated in Fig. 5. In addition, we describe the repacking strategy for improved bin utilization.
IV-C1 High-Level Search Algorithm
Algorithm 1 High-Level Search
Input: $s_{\text{high}},\textit{useRepack},\textit{requireFullPack}$
Output: $a_{\text{high}}$
1:  $v_{\text{root}}\leftarrow s_{\text{high}}$ //Initialization
2:  $\mathbb{X},\textit{solved}\leftarrow$ $\textsc{TreeExpansion}(v_{\text{root}},\emptyset,(),0,0,\textbf{false})$
3:  if $\mathbb{X}=\emptyset$ then
4:     $a_{\text{high}}\leftarrow(\texttt{terminate})$
5:     $util\leftarrow$ CurrentUtilizationRatio
6:  else
7:     $A\leftarrow$ SimulationAndGeneration($\mathbb{X}$)
8:     $\chi^{*},a_{\text{high}}\leftarrow$ BestActionSelection($\mathbb{X},A$)
9:     $util\leftarrow util(\chi^{*})$
10:  end if \* Repacking process*\
11:  if $(\mathbb{X}=\emptyset$ or NoPositionAction is in $\chi^{*})$ and useRepack then
12:     $a^{*},RepackSuccess\leftarrow$ RepackTrial($v_{\text{root}},util,\textit{requireFullPack}$)
13:     if $RepackSuccess$ then
14:        $a_{\text{high}}\leftarrow a^{*}$
15:     end if
16:  end if
17:  return  $a_{\text{high}}$
The input to Alg. 1 includes the initial high-level state $s_{\text{high}}$, a boolean flag useRepack indicating whether to attempt repacking, and a flag requireFullPack to distinguish between the following two modes: (i) Full-pack mode: This mode focuses solely on finding an optimal solution in which all known items are packed, and sub-optimal ones are not of interest. For instances where 100% bin utilization is achievable, Alg. 3 continues the search until such a solution is found. (ii) Any-time mode: In this mode, the algorithm performs progressive refinement and returns the best solution found within a limited time budget. The output of Alg. 1 is the high-level action $a_{\text{high}}$, a sequence of task primitives for bin packing.
Alg. 1 initializes the root vertex $v_{\text{root}}$ with $s_{\text{high}}$ (line 1). A tree search is then performed via the recursive function TreeExpansion (detailed in Alg. 2), which expands the search tree and returns the set of candidate packing sequences $\mathbb{X}$ along with a boolean flag solved indicating whether a solution achieving 100% bin utilization has been found (line 2). The solved remains false unless 100% utilization is reached, even if all known items can be packed.
[…]
If the initial search yields no valid placements (i.e., $\mathbb{X}=\emptyset$) or the selected sequence includes no-position actions, with repacking enabled, Alg. 1 invokes RepackTrial (Alg. 3) to iteratively unpack and re-search within the additional time (lines 1112). If repacking succeeds, the resulting action $a^{*}$ replaces $a_{\text{high}}$ (lines 1314). Finally, $a_{\text{high}}$ is returned (line 17).
IV-C2 Tree Construction
Algorithm 2 TreeExpansion
Input: $v,\mathbb{X},\chi,d,n,\textit{requireFullPack}$
Output: $\mathbb{X},\textit{solved}$
1:  $\mathcal{O}\leftarrow();$ $n=0?\ stop=\text{\bf{true}}:stop=\text{\bf{false}}$ //Initialization
2:  for each $o\in\mathcal{N}\cup\mathcal{C}$ do
3:     for each $\phi\in\{0,\pi/2\}$ do
4:        $a_{\text{low}}^{(o,\phi)}\leftarrow\pi_{\text{low}}(s_{\text{low}}^{(o,\phi)})$ and compute $r(s^{(o,\phi)}_{\text{low}},a_{\text{low}}^{(o,\phi)})$
5:        if $o$ is accessible and $a_{\text{low}}^{(o,\phi)}$ is not NoPositionAction then
[…]
7:        end if
8:        $\textsc{Append}(\mathcal{O},(o,\phi))$
9:     end for
[…]
11:        return  $\mathbb{X},\textbf{false}{}$
12:     end if
13:  end for
14:  if $stop$ then
15:     return  $\mathbb{X},\textbf{false}{}$
16:  end if
17:  $\mathcal{O}\leftarrow$ RewardSorting($\mathcal{O}$)
18:  $\mathcal{O}\leftarrow$ Selection($\mathcal{O}$)
19:  for each $(o,\phi)\in\mathcal{O}$ do
20:     $\chi^{\prime}\leftarrow\textsc{Append}(\chi,\ (o,\phi,a_{\text{low}}^{(o,\phi)},d))$
21:     if $a_{\text{low}}^{(o,\phi)}$ is not NoPositionAction then
22:        $n^{\prime}\leftarrow$ NodeCountUpdate($n,o$)
23:        $v^{\prime}\leftarrow$ ChildNodesGeneration($v,o,\phi,a_{\text{low}}^{(o,\phi)}$)
[…]
25:     end if
26:     if $a_{\text{low}}^{(o,\phi)}$ is NoPositionAction or $\mathcal{N}^{\prime}\cup\mathcal{C}^{\prime}=\emptyset$ or full then
27:        $\tilde{\chi}\leftarrow$ SequenceSorting($\chi^{\prime}$)
28:        $\mathbb{X}\leftarrow\mathbb{X}\cup\{\tilde{\chi}\}$
29:        if requireFullPack and full then
30:           return  $\mathbb{X},\textbf{true}{}$
31:        end if
32:        continue
33:     end if
34:     $\mathbb{X},\textit{solved}\leftarrow$           TreeExpansion($v^{\prime},\mathbb{X},\chi^{\prime},d+1,n^{\prime},requireFullPack$)
35:     if solved then
36:        return  $\mathbb{X},\textbf{true}{}$
37:     end if
38:  end for
39:  return  $\mathbb{X},\textbf{false}{}$
Alg. 2 recursively expands the search tree to find an optimal packing sequence for the hierarchical bin packing problem. Alg. 2 follows a DFS-BS strategy, combining depth-first expansion with selective branching. Promising branches are prioritized based on heuristic scores, and the search width is adjusted to balance exploration and computational efficiency. Alg. 2 takes as input the current state vertex $v$, the set of tuple sequences $\mathbb{X}$, the current sequence $\chi$ (which contains all tuples from the root to $v$), the current depth $d$, the number $n$ of accessible items in previous nodes, and the boolean flag requireFullPack.
Alg. 2 begins by initializing $\mathcal{O}$, which stores candidate item–orientation pairs $(o,\phi)$. The stopping flag stop is activated if no accessible items have been considered (line 1).
Each unprocessed item $o$ is considered in both orientations $\phi\in\{0,\pi/2\}$ (lines 23). For each $\phi$, Alg. 2 selects the low-level action $a_{\text{low}}^{(o,\phi)}$ using $\pi_{\text{low}}$ and evaluates its reward $r(s^{(o,\phi)}_{\text{low}},a_{\text{low}}^{(o,\phi)})$ (line 4). If an accessible item has at least one valid placement, stop is deactivated (lines 56). Considered $(o,\phi)$ pair is added to $\mathcal{O}$ (line 8). If the full-pack mode is enabled and an item $o$ yields no-position actions in both orientations, the current branch is immediately aborted (lines 1011). This avoids exploring paths that cannot lead to full utilization. If no valid placements are found for any accessible item, the current branch is aborted (lines 1415).
After evaluating all candidates, $\mathcal{O}$ is sorted by RewardSorting (line 17). This function prioritizes candidates by descending reward $r(s^{(o,\phi)}_{\text{low}},a_{\text{low}}^{(o,\phi)})$, and in case of a tie, by descending item size. Placements with higher rewards improve bin utilization, and considering larger items earlier improves overall space usage, as larger items are more difficult to fit in later stages. Then, Selection selects top-ranked candidates based on reward values (line 18). By controlling the branching factor, this strategy enhances both exploration quality and computational efficiency. Unselected candidates remain in $\mathcal{N}\cup\mathcal{C}$ and are reconsidered at deeper levels. If all candidates yield no-position actions, one is retained to preserve the current partial sequence.
Alg. 2 iterates over the sorted-selected $\mathcal{O}$ to generate child nodes (line 19). Each element of $\chi$ is a tuple of an item, its orientation, the positional action, and the node depth, as defined in (6). These values are obtained from the current recursion context and each candidate in $\mathcal{O}$. A new tuple is appended to $\chi$ for each candidate (line 20). If $a_{\text{low}}^{(o,\phi)}$ is valid, a child node $v^{\prime}$ is generated, and the count $n^{\prime}$ of accessible packable items is updated. The updated bin $B^{\prime}$ is checked for full occupancy, and the result is stored in the full flag (lines 2224). If the item cannot be placed, all items have been processed, or the bin is full, Alg. 2 sorts the current sequence $\chi^{\prime}$ into $\tilde{\chi}$ and adds it to the solution set $\mathbb{X}$ (lines 2628). The sorting prioritizes unpacked items, followed by accessible and then inaccessible ones. Among inaccessible items, arrival order determines their ordering in $\tilde{\chi}$. If both requireFullPack and full are true, Alg. 2 terminates early by returning true (lines 2930). Otherwise, since the condition in line 26 indicates no further expansion is needed, Alg. 2 skips the recursive call and proceeds to the next branch (line 32). Unless the condition in line 26 is satisfied, Alg. 2 recursively calls TreeExpansion to explore the next depth (line 34). If the recursive call returns with solved set to true, Alg. 2 terminates (lines 3536). After exploring all candidates Alg. 2 returns the set $\mathbb{X}$ (line 39).
IV-C3 Action Generation and Forward Simulation
Figure 6: Solution prioritization by total tree depth of accessible items. Among solutions yielding the same bin configuration, those that place accessible items (e.g., $o_{1}$, $o_{2}$) earlier tend to produce more compact arrangements, improving adaptability to future arrivals.
Once the tree is fully expanded, actions are generated and forward simulation is performed for each $\tilde{\chi}\in\mathbb{X}$, constructed by Alg. 2. Through forward simulation, the score $\mu(\tilde{\chi})$ is calculated from an initial value of zero. Each element of $\tilde{\chi}$ is a tuple $(o,\phi,a^{(o,\phi)}_{\text{low}},d(v^{\prime}))$. Tuples with valid positional actions are sequentially converted into pack$(o,\phi,x_{o},y_{o})$ commands. If $o$ is an unpacked or accessible item, the pack action is executed in the simulation, and the corresponding low-level reward from $a^{(o,\phi)}_{\text{low}}$ is added to $\mu(\tilde{\chi})$. In conveyor-based environments, if an item is removed from $\mathcal{N}$ via pack, subsequent items advance, potentially making previously inaccessible items accessible. If the action is a no-position action or the item remains inaccessible, the item is skipped in the simulation. The simulation terminates when either all accessible items result in no-position actions or all tuples in $\tilde{\chi}$ have been processed. Then, $util(\tilde{\chi})$, the final utilization ratio of $\tilde{\chi}$, is computed.
As the forward simulation progresses, pack actions are appended to the corresponding high-level action $a_{\tilde{\chi}}$. All pack actions for unpacked items are included. For accessible items not yet packed, at most as many pack primitives are added as the number of available robots. Other accessible items with valid positional actions are executed in the simulation but excluded from $a_{\tilde{\chi}}$ at the current step. Among accessible items, inclusion in $a_{\tilde{\chi}}$ is based on the depth value $d(v^{\prime})$. In a dual-arm setting, both robots must pick items simultaneously. The item with the smallest $d(v^{\prime})$ accessible by at least one robot is selected first. If only one robot can access it, the other is assigned the item with the next smallest $d(v^{\prime})$ it can access. If both robots can access the same item, the second smallest $d(v^{\prime})$ item accessible by either is selected next. If the next conveyor step yields no packable items, a terminate primitive is appended to $a_{\tilde{\chi}}$.
After the forward simulation, the sequence with the highest score $\mu(\tilde{\chi})$ is selected. If multiple sequences yield the same score, the one with the highest utilization ratio $util(\tilde{\chi})$ is chosen. If a tie still remains, the sequence with the smallest total $d(v^{\prime})$ for items included in the high-level action is selected, as prioritizing items with smaller depth values improves adaptability to unknown future arrivals, as illustrated in Fig. 6.
IV-C4 Repacking Solution Search
Although Alg. 2 finds a high-quality packing solution, it may not reach global optimality in the online BPP setting, where only partial item information is available at decision time. To address this, we propose an improvement procedure that enhances bin utilization by selectively unpacking and repacking items in different positions or orientations. Since Alg. 3 starts from a valid solution and can terminate at any time while progressively improving the result, it functions as an anytime algorithm.
[…]
Algorithm 3 RepackTrial
Input: $v_{\text{root}},util,\textit{requireFullPack}$
Output: $a^{*},RepackSuccess$
[…]
2:  while time limit not reached do
3:     for $i\leftarrow 1$ to $|\mathcal{I}|$ do
4:        for each subset $\mathcal{U}$ of $\mathcal{I}$ with $i$ elements (last-packed-first) do
5:           Clone $B,\mathcal{N},\mathcal{I},\mathcal{C}$ to $\hat{B},\hat{\mathcal{N}},\hat{\mathcal{I}},\hat{\mathcal{C}}$
6:           for each item $o$ in $\mathcal{U}$ do
7:              Execute $\texttt{unpack}(o)$ and update $\hat{v}_{\text{root}}$ with $\hat{B},\hat{\mathcal{N}},\hat{\mathcal{I}},\hat{\mathcal{C}}$
8:              $a\leftarrow\textsc{Append}(a,\texttt{unpack}(o))$
9:           end for
10:           $\mathbb{X},\textit{solved}\leftarrow$              $\textsc{TreeExpansion}(\hat{v}_{\text{root}},\emptyset,(),0,0,\textit{requireFullPack})$
11:           if $\mathbb{X}\neq\emptyset$ then
12:              $A\leftarrow$ SimulationAndGeneration($\mathbb{X}$)
13:              $\chi^{*},a_{\chi^{*}}\leftarrow$ BestActionSelection($\mathbb{X},A$)
14:              $a_{\chi^{*}}\leftarrow\textsc{ActionUpdate}(a,a_{\chi^{*}})$
15:              if requireFullPack and NoPositionAction is not in $\chi^{*}$ then
[…]
17:                 return  $a_{\chi^{*}},RepackSuccess$
18:              else if not requireFullPack and $util<util(\chi^{*})$ then
[…]
20:                 $a^{*}\leftarrow a_{\chi^{*}},util\leftarrow util(\chi^{*})$
21:              end if
22:           end if
23:           Reset $a\leftarrow()$
24:        end for
25:     end for
26:  end while
27:  return  $a^{*},RepackSuccess$
IV-D Task Allocation and Atomic Action Sequencing
Until now, action primitives have been generated without being assigned to specific robots. While single-arm systems have no ambiguity, dual-arm systems require explicit task distribution. pack primitives for newly arrived items are allocated to the robot closest in Euclidean distance. The primitives unpack are assigned alternately, starting with the less-loaded robot. pack for unpacked items is assigned to the same robot that executed the corresponding unpack, with repacking done in reverse order of unpacking. Robots resume packing new items after completing all unpacking and repacking tasks.
To generate feasible motions for executing the assigned task primitives, we decompose them into atomic actions. As dual-arm motion planning in a shared task space is an inherently complex problem, we sequence actions in a synchronized manner rather than planning full asynchronous motions. This strategy reduces planning time at the expense of execution time. Our atomic action sequencing guarantees that robots do not operate in overlapping task space regions simultaneously.
In our environment, we define three task subspaces: the bin, the conveyor belt, and the buffer space of each robot. Both pack and unpack are decomposed into two atomic actions: pick and place to (location).
  • pack: pick from the conveyor or buffer, then place to bin.
  • unpack: pick from the bin, then place to buffer.
We also define two additional atomic actions: standby, where a robot waits, and ready, where it returns to its initial pose after completing all assigned tasks.
Atomic action sequencing is constructed as follows:
  1. 1.
    Assign standby to the robot with fewer tasks.
  2. 2.
    Add pick and place to buffer for all unpack from the first without a corresponding pack to the last with one.
  3. 3.
    If the last unpack is followed by a corresponding pack, omit place to buffer and pick, yielding pickplace to bin.
  4. 4.
    For all remaining pack, add pick and place to bin.
  5. 5.
    Conclude with a ready action for each robot.
Adding standby to one robot ensures that the two tobots do not operate in overlapping regions simultaneously, simplifying dual-arm motion planning. Any off-the-shelf motion planner can be used; we adopt an efficient variant suitable for multi-arm coordination.
TABLE I: Scenario configurations with varying robot and item setups.
Scenario
S-R1
A1
S-R5
A1
S-R5
A3
D-R2
A2O2
D-R5
A2O2
D-R5
A3O1
Fig. 1 (a) (b) (c) (d) (e) (f)
Num. of Robots 1 2
Max Num. of
Recognized Items
1 5 5 2 5 5
Num. of Accessible
Items per Robot
1 1 3 2 2 3
Num. of Items
in Overlapping
Workspace
- - - 2 2 1
V Experiments
We evaluate our hierarchical bin packing framework using a hybrid architecture, integrating task planning, DRL, and motion planning modules across two systems. The algorithm runs on a machine with an AMD Ryzen 7 9700X CPU, 64 GB RAM, and a GeForce RTX 4080 SUPER GPU. Simulation and motion planning are executed on a separate system with an AMD Ryzen 7 5800X CPU, 32 GB RAM, and a GeForce RTX 3070 Ti GPU. Modules communicate via ROS1 [26] and TCP/IP to synchronize the interaction between the planner and simulation. The DRL agent is implemented in PyTorch [27], and the simulation environment is developed in Unity [28], providing physical dynamics. All training and evaluation experiments are conducted in a $10\times 10$ bin environment.
V-A Dataset
We construct two types of datasets: Random instances, used only for testing, and 100% set instances, used for both training and testing.
In the Random instance, items are generated by sampling the width and height of each item uniformly, with each dimension limited to at most half the bin size. A large number of sets are included to ensure sufficient diversity and test coverage.
The 100% set instances are constructed to train the RL agent and evaluate the algorithm. It consists of item sets where full bin utilization (100%) is achievable. These sets enable the agent to learn optimal packing strategies and serve as a benchmark to test whether repacking can fully utilize the bin space. To construct each set, item dimensions are randomly sampled and placed at feasible locations to ensure compact, valid placements. Oversized items are avoided by considering remaining space during sampling. Once the bin is fully packed, the item order is shuffled to simulate online arrival. In total, 9.2 million item sets are generated across 100% set instances: 7.2 million (78%) are used to train the A3C worker networks, and 2 million (22%) are reserved exclusively for evaluating the global network.
V-B Training Results
(a) Policy Loss
(b) Value Loss
(c) Reward
Figure 7: Training curves for the proposed RL agent. (a) Policy loss, (b) value loss, and (c) reward.
Our A3C agent uses one global network and three parallel worker networks. Each worker independently interacts with its own environment and periodically updates the global network. This architecture enables stable learning by leveraging diverse experiences and reducing correlation among samples. The agent is implemented in PyTorch [27] and trained for 7.2 million episodes (2.4 million per worker), requiring approximately 28 hours. Fig. 7 shows the training progress of the agent. The policy loss decreases over time, indicating that the agent improves its action selection. The value loss reflects the performance of the critic in predicting state values and also shows a downward trend, demonstrating stable learning of the value function. The reward, obtained per episode, increases steadily throughout the training process, confirming the effectiveness of the learned policy in producing compact and efficient packing results. These results validate that the A3C-based agent can learn a meaningful low-level policy capable of generalizing to diverse packing situations, serving as a reliable subroutine for the high-level search algorithm.
V-C Algorithm Tests
V-C1 Scenarios
TABLE II: Packing performance across different scenarios.
Scenario S-R1A1 S-R5A1 S-R5A3 D-R2A2O2 D-R5A2O2 D-R5A3O1
Rotation No Yes No Yes No Yes No Yes No Yes No Yes
Mean Utilization (%) 78.88 85.17 86.45 92.38 94.77 98.21 85.94 91.57 91.99 96.46 95.86 98.59
Total Packed Items 8715 9422 9559 10229 10488 10863 9507 10127 10179 10661 10602 10905
Total Packing Steps Same as total packed items 5438 5706 6128 6333 5787 5893
(a) Performance on Random instances.
Scenario S-R1A1 S-R5A1 S-R5A3 D-R2A2O2 D-R5A2O2 D-R5A3O1
Repacking No Yes No Yes No Yes No Yes No Yes No Yes
Mean Utilization (%) 81.52 100 95.54 100 97.67 100 86.98 100 96.49 100 96.53 100
Num. of Full Bins 684 3000 2213 3000 2395 3000 865 3000 1999 3000 2240 3000
Mean Num. of
Packed Items
6.85 8.17 7.85 8.17 7.97 8.17 7.44 8.17 7.84 8.17 7.91 8.17
Mean Num. of
Packing Steps
Same as mean of packed items 4.16 4.34 4.40 4.34 4.23 4.34
Mean Num. of
Repacking Items
- 2.42 - 0.39 - 0.30 - 1.90 - 0.52 - 0.39
Mean Num. of Steps
including Repacking
- 1.00 - 0.27 - 0.21 - 0.80 - 0.34 - 0.26
Mean Taken Time (s)
for Repacking Solution
- 2.55 - 1.43 - 0.90 - 2.82 - 1.33 - 0.79
(b) Performance on 100% set with and without repacking.
We define six distinct scenarios to evaluate our bin packing framework under varying robotic configurations and perception constraints, as summarized in Table I and shown in Fig. 1. Each scenario label encodes the following:
  • S (single robot) or D (dual robot): robot configuration,
  • R: maximum number of recognized items,
  • A: number of accessible items per robot,
  • O: number of items in the overlapping workspace (for dual-robot settings only).
Scenarios S-R1A1, S-R5A1, and S-R5A3 represent single-arm configurations, while D-R2A2O2, D-R5A2O2, and D-R5A3O1 correspond to dual-arm settings. S-R1A1 denotes a minimal setup where one robot can recognize and access only a single item. In contrast, D-R5A3O1 reflects the most capable configuration, where two robots can recognize five items, each access three, and share one item in the overlapping region.
V-C2 Test with Random Instances
We evaluate our method on the Random instances with 1000 episodes per scenario, where each episode involves packing a single bin. All scenarios share the same item arrival sequence across the entire test. We first evaluate our method without repacking across six scenarios, with and without item rotation. As shown in Table II(a), both bin utilization and the total number of packed items consistently improve when rotation is enabled. For instance, in Scenario S-R1A1, the utilization increases from 78.88% to 85.17%, and the number of packed items from 8715 to 9422. Similar trends are observed across all scenarios, with Scenario D-R5A3O1 achieving the highest utilization of 98.59%. We then test repacking in four scenarios where there is room for improvement: S-R1A1, S-R5A1, D-R2A2O2, and D-R5A2O2. Fig. 8 shows that even a short planning time of 1 second for repacking significantly improves the final bin utilization. In Scenario S-R1A1, the utilization increases from 85.17% to 98.01%, and in D-R2A2O2 from 91.57% to 99.03%. In Scenario S-R5A1, the utilization improves from 92.38% to 99.20%, and in D-R5A2O2 from 96.46% to 99.67%. Scenarios with more known items result in fewer repacking actions, which helps reduce motion overhead during execution.
(a) Bin utilization over time
(b) Total number of repacked items
Figure 8: Impact of repacking planning time on packing performance. (a) Bin utilization improves as more planning time is allocated. (b) The total number of repacked items increases with longer planning time.
V-C3 Test with 100% set
We evaluate our method on the 100% set with 3000 episodes per scenario, where each episode uses the same fixed item arrival sequence. As shown in Table II(b), our method achieves high utilization even without repacking. When repacking is enabled, 100% utilization is achieved in all scenarios. The number of fully packed bins increases significantly with repacking, from 684 to 3000 in S-R1A1 and from 865 to 3000 in D-R2A2O2. Repacking improves packing quality at the cost of moderate computation time, with an average of less than 3 seconds to find a repacking solution.
V-C4 Compare with Baseline
TABLE III: Comparison of bin utilization (%) with baseline methods in $4\times 4$ and $5\times 5$ bin environments.
Method Uti. (%)
Deep-Pack 94.87
BERM 97.23
Ours 98.23
(a) $4\times 4$ bin
Method Uti. (%)
ShelfNextFit 75.80
Skyline 85.33
Deep-Pack 91.48
BERM 94.30
Ours 96.58
(b) $5\times 5$ bin
To validate the effectiveness of our method, we compare it with existing baselines, including heuristic methods (Skyline [4], ShelfNextFit [5]) and learning-based approaches (Deep-Pack [7], BERM [19]). The evaluation follows the same setup as the baselines in $4\times 4$ and $5\times 5$ bin environments under a fully online setting without rotation or repacking, ensuring fair comparison. Table III shows that our method achieves the highest mean bin utilization in both settings—98.23% in the $4\times 4$ bin (compared to 94.87% for Deep-Pack and 97.23% for BERM), and 96.58% in the $5\times 5$ bin (compared to 91.48% and 94.30%, respectively). These results demonstrate that our DRL-based policy outperforms the state-of-the-art methods in a constrained online scenario.
V-D Simulation Tests
The simulation environment is developed using Unity [28] to evaluate the proposed framework under physical constraints. Universal Robots UR5e is used as the manipulator in both single- and dual-arm configurations. The simulation environments for each setting are shown in Fig. 9. For motion planning, cuRobo [29] is employed in both configurations, while StopNGo [13] is integrated to coordinate dual-arm execution. Each setup is tested over 30 episodes using the 100% set instances.
As shown in Table IV, the dual-arm scenario achieves a shorter makespan than the single-arm scenario, despite requiring longer per-step planning durations. Specifically, it reduces the overall makespan by approximately 24% in both repacking and non-repacking cases. Since planning and execution proceed concurrently, planning for the next step is typically completed during the current execution, minimizing idle time. Although repacking introduces slight overhead in planning and execution, it improves bin utilization (as discussed in Sec. V-C3), which justifies the additional cost. In instances with an odd number of items, the final item is always handled by a single robot, limiting the relative gain in makespan reduction. These results demonstrate that the proposed framework can generate physically feasible and time-efficient plans for dual-arm robotic systems.
(a) S-R5A3 simulation environment
(b) D-R5A3O1 simulation environment
Figure 9: Simulation environments used for testing. (a) single-arm setting (S-R5A3), and (b) dual-arm setting (D-R5A3O1).
TABLE IV: Planning and Execution Time in Single- and Dual-Arms
Scenario S-R5A3 D-R5A3O1
Repacking No Yes No Yes
Task Planning Time (s) 0.71 0.76 0.77 0.85
Motion Planning Time (s) 0.73 0.77 1.46 1.60
Execution Time (s) 15.50 15.93 21.22 22.39
Average Makespan
per episode (s)
124.48 130.98 94.29 100.12
VI Conclusion
This paper presented a hierarchical bin packing framework that combines heuristic tree search and deep reinforcement learning to solve 2D bin packing problems. The method supports repacking and dual-arm coordination through integrated task and motion planning. A low-level RL agent proposes placement positions, while the high-level search explores packing and unpacking strategies under item availability and accessibility constraints. Experiments demonstrate near-optimal bin utilization and show that the framework generates physically feasible, time-efficient plans, reducing execution time in dual-arm scenarios. These results highlight the benefits of repacking and parallel execution for practical robotic systems.

### v2

@@
 It was tested in simulation using UR5e robots and open-source planning tools, demonstrating effective use of bin space and faster task execution with the dual-arm setup compared to the single-arm setup.
-Index Terms:
-Bin packing, dual manipulators, warehouse automation, hierarchical planning
 I Introduction
@@
 Semi-online BPPs are more common in industrial contexts such as logistics and manufacturing.
+In typical warehouse or inspection setups, a camera or barcode reader is installed above a conveyor belt, allowing partial observation of incoming items within a fixed field of view.
+In such cases, the system can access limited item information (e.g., size, orientation, or label data) before packing, which aligns with the assumptions of the 2D online and semi-online settings.
 In such settings, strategies for managing uncertainty, such as repacking, can improve adaptability to changes in item arrivals.
 BPP is also categorized by packing dimension (e.g., 2D, 3D) [12].
-Although there has been a line of research on both 2D BPP and 3D BPP, 2D BPP remains underexplored, particularly in achieving near-optimal packing through rotation and/or repacking.
+Although there has been a line of research on both 2D BPP and 3D BPP, many real-world applications are effectively 2D due to physical or operational constraints.
+For example, fragile or breakable items such as glass panels and ceramics cannot be stacked; inspection or vision-based operations often require items to remain within a single camera view; and layout optimization on flat materials such as textiles, leather, or printed circuit boards (PCBs) is inherently 2D.
+In addition, certain industrial containers, such as interlocking plastic crates (e.g., milk crates), and uniformly shaped items that can be stacked in shallow multi-layer configurations, also fit naturally within a 2D formulation.
+In these environments, the packing task is dominated by footprint efficiency rather than vertical stacking.
+Motivated by these practical use cases, this work focuses on the 2D online and semi-online bin packing problem.
 Most existing BPP methods assume a single-robot setting, although coordinating multiple manipulators can significantly reduce execution time through parallelization.
 Coordinating multiple robots for BPP remains computationally challenging due to the inherent complexity of multi-robot task allocation and multi-arm motion planning [13, 14, 15, 16].
+While asynchronous or semi-synchronous operation can potentially improve execution efficiency, these approaches considerably increase the complexity and computation time of task and motion planning required to ensure safe and collision-free coordination.
+Therefore, the proposed framework adopts a synchronous operation, which provides predictable coordination between manipulators, simplifies collision avoidance, and reduces planning complexity, which, makes it more practical for real-world deployment.
 This paper addresses the 2D bin packing problem (BPP) under practical scenarios involving varying numbers of known items, accessible items, and manipulators, as illustrated in Fig.
@@
   • A DRL framework that outperforms the state-of-the-art 2D online BPP methods with significantly higher bin utilization.
-  • A heuristic search algorithm that supports a wide range of 2D BPP scenarios, including offline, semi-online, and online settings, with varying levels of item accessibility and information, and even multi-manipulator configurations.
+  • A heuristic search algorithm that integrates robot accessibility and multi-manipulator coordination into the policy generation process, enabling feasible and optimized packing decisions under offline, semi-online, and online conditions with varying perception ranges.
   • A repacking strategy that consistently achieves near-100% bin utilization.
@@
 We also discuss approaches for coordinating multi-robot operations in shared workspaces.
+Table I summarizes representative 2D and 3D BPP studies, focusing on their problem settings and robotic implementation aspects such as repacking, multi-bin handling, and dual-arm coordination.
+TABLE I: Comparison of representative studies in 2D and 3D bin packing settings.
+Characteristic 2D 3D
+Offline / Semi-online Ours, [3] [6, 18, 8, 19, 20, 21, 22, 23, 24, 25]
+Online Ours,[7, 26, 4, 5] [19, 20, 21, 22, 23]
+Repacking Ours [22, 23, 24]
+Multi-bin [6, 18, 19, 24]
+Robotic Implementation Ours [18, 19, 20, 22, 23, 24, 25]
+Dual-arm Allocation per bin Ours
 In the offline BPP where all items are known, metaheuristic approaches such as genetic algorithms [3], Tabu Search [6], and layer-building strategies [18] have been proposed.
@@
 Additionally, [22] compares unpacked and packed states unfairly, since the unpacked state has fewer packed items and thus an advantage in value.
+Recent studies have also explored multi-arm in bin packing scenarios.
+[24] proposes a deep reinforcement learning framework for online and concurrent 3D bin packing with bin replacement, in which multiple bins are processed in parallel to improve throughput.
+Their work focuses on optimizing packing efficiency across multiple containers, rather than coordinating multiple manipulators within a single bin.
+On the other hand,  [25] develops a dual-arm robotic packing framework emphasizing joint motion coordination and trajectory planning for cooperative manipulation of a single object.
+While both studies address related challenges in robotic packing, their problem formulations differ from ours: we focus on dual-manipulator cooperation for packing multiple objects within a single bin under online and semi-online settings, where task allocation and placement decisions are jointly optimized through a hierarchical algorithm.
 While multi-robot manipulation improves packing throughput, it introduces coordination challenges.
 Some works have addressed item rearrangement in shared workspaces [14, 15, 16], but coordinated bin packing remains largely unexplored.
-To fill this gap, we propose a framework for 2D BPP under realistic multi-robot settings.
-Our method considers item accessibility through forward simulation to ensure feasibility, and integrates coordinated task planning and repacking within a unified hierarchical structure.
-Leveraging a dual-manipulator system, we enable parallel execution while ensuring packing decisions are both feasible and optimized for bin utilization.
+To fill this gap, we propose a framework for 2D BPP under realistic dual-arm settings.
+Unlike previous studies that determine object placement and orientation before considering robotic execution, our method jointly plans the packing policy while accounting for the reachability of both single and dual manipulators.
+This integrated formulation ensures that packing decisions are physically feasible for robot execution while maintaining near-optimal packing efficiency.
+Moreover, the framework synergistically integrates a low-level RL policy with a high-level tree search, combining the flexibility of learning-based decision making with the reliability of search-based reasoning.
 III Problem Formulation
-(a) The workspace and cache of the dual manipulators
+(a) The task space and cache of the dual manipulators
 (b) The bin and an item
@@
 Figure 2: Illustration of the Bin Packing system.
-(a) The manipulator is denoted by $m$, $\mathcal{Z}$ represents the workspace of the manipulator, and $C$ is the temporary storage area for unpacked items.
+(a) The manipulator is denoted by $m$, $\mathcal{Z}$ represents the task space of the manipulator, and $C$ is the temporary storage area for unpacked items.
 (b) The bin is an open cuboid, and the next item $o$ is a solid cuboid.
-(c) The bin is modeled as a padded binary grid matrix $B$, where each cell corresponds to a low-level position action indexed by its coordinate in the image coordinate system.
+(c) The bin is modeled as a padded binary grid map $B$, where each cell corresponds to a low-level position action indexed by its coordinate in the image coordinate system.
 The values increase row-wise from left to right and top to bottom, with the origin located at the top-left corner.
@@
 2(a) and 2(b), we consider a 2D BPP environment where the bin is modeled as an open cuboid and items as closed cuboids, all projected onto the $xy$-plane.
-We assume up to two robotic manipulators, denoted by $m_{i}$ for $i\in\{1,2\}$, each with a reachable workspace $\mathcal{Z}_{i}\subset\mathcal{W}\subseteq\mathbb{R}^{3}$.
+We assume up to two robotic manipulators, denoted by $m_{i}$ for $i\in\{1,2\}$, each with a task space $\mathcal{Z}_{i}\subset\mathcal{W}\subseteq\mathbb{R}^{3}$.
 Items arrive sequentially via a conveyor belt (or other similar facilities), which advances as items are removed.
 The bin and each item $o$ are discretized into grid cells: the bin is a $W\times H$ grid, and item $o$ occupies $w_{o}\times h_{o}$ cells, where $W,H,w_{o},h_{o}\in\mathbb{Z}^{+}$.
-The bin occupancy is represented by a padded binary matrix $B\in\{0,1\}^{(W+2)\times(H+2)}$, where 00 and $1$ indicate free and occupied cells.
+The bin occupancy is represented by a padded binary map $B\in\{0,1\}^{(W+2)\times(H+2)}$, where 0 and $1$ indicate free and occupied cells.
 The outermost rows and columns are padded with $1$s to represent the bin boundaries (see Fig.
@@
 Each manipulator $m_{i}$ has a buffer space $C_{i}\subset\mathcal{Z}_{i}$ to temporarily store items in $\mathcal{C}$.
-The High-level MDP
+III-A The High-level MDP
 The high-level MDP governs overall task planning, which is defined as $(\mathcal{S}_{\text{high}},\mathcal{A}_{\text{high}},P_{\text{high}},R_{\text{high}})$.
@@
 $\texttt{pack}(o_{1},\phi_{o_{1}},x_{o_{1}},y_{o_{1}}),\texttt{pack}(o_{2},\phi_{o_{2}},x_{o_{2}},y_{o_{2}}),$
-[…]
+$\texttt{terminate})$
 unpacks $o_{1}$ and $o_{2}$, then repacks them at new positions, followed by termination. An action may consist solely of terminate, or include only pack.
@@
 The cumulative high-level reward indicates the bin utilization ratio, encouraging space-efficient packing in real-world applications.
-The Low-level MDP
+III-B The Low-level MDP
 As mentioned, the primitive $\texttt{pack}(o,\phi_{o},x_{o},y_{o})$ requires a specific position $(x_{o},y_{o})$ to place item $o$ in the bin.
@@
 The high-level MDP evaluates both orientations and selects one along with its corresponding position to complete the pack primitive.
-The bin position is represented by an index from 00 (top-left) to $W\cdot H-1$ (buttom-right), following a row-wise order as illustrated in Fig.
+The bin position is represented by an index from 0 (top-left) to $W\cdot H-1$ (buttom-right), following a row-wise order as illustrated in Fig.
 2(c).
@@
 $a_{\text{low}}^{(o,\phi)}=\pi_{\text{low}}(s_{\text{low}}^{(o,\phi)})=\pi_{\text{low}}(\{B,\mathbf{l}_{o}(\phi)\}).$ (5)
-This leads to a new vertex $v^{\prime}=\{B^{\prime},\mathcal{N}^{\prime},\mathcal{I}^{\prime},\mathcal{C}^{\prime}\}\in\mathcal{V}$ in the tree, where $B^{\prime}$ is updated via (1), and item $o$ is moved from $\mathcal{N}$ or $\mathcal{C}$ to $\mathcal{I}$ [^1] Item $o$ is removed from $\mathcal{N}$ or $\mathcal{C}$, whichever contains it, and added to $\mathcal{I}$.
+This leads to a new vertex $v^{\prime}=\{B^{\prime},\mathcal{N}^{\prime},\mathcal{I}^{\prime},\mathcal{C}^{\prime}\}\in\mathcal{V}$ in the tree, where $B^{\prime}$ is updated via (1), and item $o$ is moved from $\mathcal{N}$ or $\mathcal{C}$ to $\mathcal{I}$.
+[^1] Item $o$ is removed from $\mathcal{N}$ or $\mathcal{C}$, whichever contains it, and added to $\mathcal{I}$.
 If $a_{\text{low}}^{(o,\phi)}$ is the no-position action (2), no child is generated and $v$ becomes a leaf vertex.
@@
 Output: $a_{\text{high}}$
-1:  $v_{\text{root}}\leftarrow s_{\text{high}}$ //Initialization
+1:  $v_{\text{root}}\leftarrow s_{\text{high}}$ .//Initialization
 2:  $\mathbb{X},\textit{solved}\leftarrow$ $\textsc{TreeExpansion}(v_{\text{root}},\emptyset,(),0,0,\textbf{false})$
@@
 12:   $a^{*},RepackSuccess\leftarrow$ RepackTrial($v_{\text{root}},util,\textit{requireFullPack}$)
-13:     if $RepackSuccess$ then
+13:   if RepackSuccessRepackSuccess then
 14:    $a_{\text{high}}\leftarrow a^{*}$
@@
 The solved remains false unless 100% utilization is reached, even if all known items can be packed.
-[…]
+If no valid solution is found ($\mathbb{X}=\emptyset$), Alg.
+1 sets $a_{\text{high}}$ to terminate to end the episode, and assigns utilutil, the highest achievable utilization ratio identified so far, as the current value (lines 35).
+Otherwise, each sequence $\tilde{\chi}\in\mathbb{X}$ is evaluated, and the candidate high-level action set $A$ is generated through SimulationAndGeneration (line 7).
+Then, BestActionSelection selects the best sequence $\chi^{*}$ and sets its action as $a_{\text{high}}$ (line 8).
+Both procedures are detailed in Sec.
+IV-C3.
+After $\chi^{*}$ is determined, utilutil is assigned as $\mathit{util}(\chi^{*})$, the utilization ratio computed from $\chi^{*}$ (line 9).
 If the initial search yields no valid placements (i.e., $\mathbb{X}=\emptyset$) or the selected sequence includes no-position actions, with repacking enabled, Alg.
@@
 Output: $\mathbb{X},\textit{solved}$
-1:  $\mathcal{O}\leftarrow();$ $n=0?\ stop=\text{\bf{true}}:stop=\text{\bf{false}}$ //Initialization
+1:  $\mathcal{O}\leftarrow();$ $n=0?\ stop=\text{\bf{true}}:stop=\text{\bf{false}}$ .//Initialization
 2:  for each $o\in\mathcal{N}\cup\mathcal{C}$ do
@@
 5:    if $o$ is accessible and $a_{\text{low}}^{(o,\phi)}$ is not NoPositionAction then
-[…]
+6:     $\textit{stop}\leftarrow\textbf{false}{}$
 7:    end if
@@
 9:   end for
-[…]
+10:   if requireFullPack and $(a_{\text{low}}^{(o,0)},a_{\text{low}}^{(o,\pi/2)}\text{ are NoPositionAction})$ then
 11:    return $\mathbb{X},\textbf{false}{}$
@@
 13:  end for
-14:  if $stop$ then
+14:  if stopstop then
 15:   return $\mathbb{X},\textbf{false}{}$
@@
 22:    $n^{\prime}\leftarrow$ NodeCountUpdate($n,o$)
-23:        $v^{\prime}\leftarrow$ ChildNodesGeneration($v,o,\phi,a_{\text{low}}^{(o,\phi)}$)
 […]
+24:    $\textit{full}\leftarrow$ BinFullCheck($B^{\prime}$)
 25:   end if
@@
 Among solutions yielding the same bin configuration, those that place accessible items (e.g., $o_{1}$, $o_{2}$) earlier tend to produce more compact arrangements, improving adaptability to future arrivals.
-Once the tree is fully expanded, actions are generated and forward simulation is performed for each $\tilde{\chi}\in\mathbb{X}$, constructed by Alg.
-2.
-Through forward simulation, the score $\mu(\tilde{\chi})$ is calculated from an initial value of zero.
-Each element of $\tilde{\chi}$ is a tuple $(o,\phi,a^{(o,\phi)}_{\text{low}},d(v^{\prime}))$.
-Tuples with valid positional actions are sequentially converted into pack$(o,\phi,x_{o},y_{o})$ commands.
-If $o$ is an unpacked or accessible item, the pack action is executed in the simulation, and the corresponding low-level reward from $a^{(o,\phi)}_{\text{low}}$ is added to $\mu(\tilde{\chi})$.
-In conveyor-based environments, if an item is removed from $\mathcal{N}$ via pack, subsequent items advance, potentially making previously inaccessible items accessible.
-If the action is a no-position action or the item remains inaccessible, the item is skipped in the simulation.
-The simulation terminates when either all accessible items result in no-position actions or all tuples in $\tilde{\chi}$ have been processed.
-Then, $util(\tilde{\chi})$, the final utilization ratio of $\tilde{\chi}$, is computed.
+[…]
 As the forward simulation progresses, pack actions are appended to the corresponding high-level action $a_{\tilde{\chi}}$.
@@
 Among accessible items, inclusion in $a_{\tilde{\chi}}$ is based on the depth value $d(v^{\prime})$.
-In a dual-arm setting, both robots must pick items simultaneously.
+In the proposed synchronized dual-arm setting, two objects are selected and assigned to the respective robots.
 The item with the smallest $d(v^{\prime})$ accessible by at least one robot is selected first.
@@
 If the next conveyor step yields no packable items, a terminate primitive is appended to $a_{\tilde{\chi}}$.
-After the forward simulation, the sequence with the highest score $\mu(\tilde{\chi})$ is selected.
-If multiple sequences yield the same score, the one with the highest utilization ratio $util(\tilde{\chi})$ is chosen.
-If a tie still remains, the sequence with the smallest total $d(v^{\prime})$ for items included in the high-level action is selected, as prioritizing items with smaller depth values improves adaptability to unknown future arrivals, as illustrated in Fig.
-6.
+[…]
 IV-C4 Repacking Solution Search
@@
 Output: $a^{*},RepackSuccess$
-[…]
+1:  $a^{*}\leftarrow(),a\leftarrow(),RepackSuccess$ $\leftarrow\textbf{false}{}$ .//Initialization
 2:  while time limit not reached do
@@
 15:      if requireFullPack and NoPositionAction is not in $\chi^{*}$ then
-[…]
+16:       RepackSuccessRepackSuccess $\leftarrow\textbf{true}{}$
 17:       return $a_{\chi^{*}},RepackSuccess$
 18:      else if not requireFullPack and $util<util(\chi^{*})$ then
-[…]
+19:       RepackSuccessRepackSuccess $\leftarrow\textbf{true}{}$
 20:       $a^{*}\leftarrow a_{\chi^{*}},util\leftarrow util(\chi^{*})$
@@
 Atomic action sequencing is constructed as follows:
-  1. 1.
+  1. (i)
     Assign standby to the robot with fewer tasks.
-  2. 2.
+  2. (ii)
     Add pick and place to buffer for all unpack from the first without a corresponding pack to the last with one.
-  3. 3.
+  3. (iii)
     If the last unpack is followed by a corresponding pack, omit place to buffer and pick, yielding pickplace to bin.
-  4. 4.
+  4. (iv)
     For all remaining pack, add pick and place to bin.
-  5. 5.
+  5. (v)
     Conclude with a ready action for each robot.
@@
 Any off-the-shelf motion planner can be used; we adopt an efficient variant suitable for multi-arm coordination.
-TABLE I: Scenario configurations with varying robot and item setups.
+TABLE II: Scenario configurations with varying robot and item setups.
 Scenario
@@
 All training and evaluation experiments are conducted in a $10\times 10$ bin environment.
+Considering the placement margin, robot positioning accuracy, and camera perception error, the 10 × 10 grid resolution is sufficiently precise for stable packing and is therefore adopted as the representative configuration for experimental validation.
 V-A Dataset
 We construct two types of datasets: Random instances, used only for testing, and 100% set instances, used for both training and testing.
-In the Random instance, items are generated by sampling the width and height of each item uniformly, with each dimension limited to at most half the bin size.
-A large number of sets are included to ensure sufficient diversity and test coverage.
-The 100% set instances are constructed to train the RL agent and evaluate the algorithm.
-It consists of item sets where full bin utilization (100%) is achievable.
-These sets enable the agent to learn optimal packing strategies and serve as a benchmark to test whether repacking can fully utilize the bin space.
-To construct each set, item dimensions are randomly sampled and placed at feasible locations to ensure compact, valid placements.
-Oversized items are avoided by considering remaining space during sampling.
-Once the bin is fully packed, the item order is shuffled to simulate online arrival.
-In total, 9.2 million item sets are generated across 100% set instances: 7.2 million (78%) are used to train the A3C worker networks, and 2 million (22%) are reserved exclusively for evaluating the global network.
+In the Random instance, items are generated by independently sampling each dimension uniformly: $w_{o}\sim\mathcal{U}(1,\lfloor W/2\rfloor)$ and $h_{o}\sim\mathcal{U}(1,\lfloor H/2\rfloor)$, ensuring no single item exceeds half the bin size in either dimension.
+100% Set instances consist of item sets where full bin utilization (100%) is achievable, enabling the agent to learn optimal packing strategies and serving as a benchmark to test whether repacking can fully utilize the bin space.
+The construction procedure is detailed in Alg.
+4.
+Item dimensions are sampled from Gaussian-weighted probabilities (standard deviation $\sigma=2$) over width and height, which reduces the probability of generating items with dimension 1 in early stages.
+This design prevents an overabundance of such small or elongated items that would inevitably arise later when filling remaining gaps.
+For each sampled item, we compute edge-contact rewards for all feasible placement positions and greedily select the position that maximizes this reward, with random tie-breaking when multiple positions share the maximum value.
+Oversized items that cannot fit in the remaining space are rejected and resampled.
+Once the bin is full or only one cell remains empty, the process is terminated.
+In the case where one cell remains empty, a $1\times 1$ item is added to achieve exact 100% fill.
+Finally, the item order is shuffled to simulate online arrival.
+In total, 9.2 million item sets are generated: 7.2 million (78%) are used to train the A3C worker networks, and 2 million (22%) are reserved exclusively for evaluating the global network.
+Algorithm 4 100% Set Instance Generation
+0:  Bin size $W\times H$, Gaussian std $\sigma=2$
+0:  Item set $\mathcal{S}$ with 100% bin utilization
+1:  $\mathcal{S}\leftarrow\emptyset$, $\text{area}\leftarrow 0$
+2:  Initialize empty bin $B$ of size $W\times H$
+3:  $\max_{w}\leftarrow 0$, $\max_{h}\leftarrow 0$
+4:  $p_{w}\leftarrow\text{GaussianProb}(W,\sigma)$
+5:  $p_{h}\leftarrow\text{GaussianProb}(H,\sigma)$
+6:  while $\text{area}<W\times H-1$ do
+7:   repeat
+8:    repeat
+9:     Sample $w_{o}\sim p_{w}$, $h_{o}\sim p_{h}$
+10:    until $(w_{o}+h_{o}\leq W+H-\max_{w}-\max_{h})$ AND    $(\text{area}+w_{o}\cdot h_{o}\leq W\times H)$
+11:    Initialize $R[y,x]\leftarrow$-$\infty$ for all $(x,y)$
+12:    Compute $R[y,x]\leftarrow$ $\text{EdgeContact}(B,x,y,w_{o},h_{o})$ for feasible $(x,y)$
+13:   until $\max_{(x,y)}R[y,x]\geq 0$
+14:   $(x^{*},y^{*})\leftarrow\arg\max_{(x,y)}R[y,x]$ with random tie-breaking
+15:   Place item $(w_{o},h_{o})$ at position $(x^{*},y^{*})$ in bin $B$
+16:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{(w_{o},h_{o})\}$
+17:   $\text{area}\leftarrow\text{area}+w_{o}\cdot h_{o}$
+18:   $\max_{w}\leftarrow\max(\max_{w},w_{o})$, $\max_{h}\leftarrow\max(\max_{h},h_{o})$
+19:  end while
+20:  if $\text{area}=W\times H-1$ then
+21:   $\mathcal{S}\leftarrow\mathcal{S}\cup\{(1,1)\}$
+22:  end if
+23:  Randomly shuffle $\mathcal{S}$
+24:  return $\mathcal{S}$
 V-B Training Results
@@
 These results validate that the A3C-based agent can learn a meaningful low-level policy capable of generalizing to diverse packing situations, serving as a reliable subroutine for the high-level search algorithm.
-V-C Algorithm Tests
-V-C1 Scenarios
-TABLE II: Packing performance across different scenarios.
+TABLE III: Packing performance across different scenarios.
 Scenario S-R1A1 S-R5A1 S-R5A3 D-R2A2O2 D-R5A2O2 D-R5A3O1
@@
 Total Packing Steps Same as total packed items 5438 5706 6128 6333 5787 5893
-(a) Performance on Random instances.
+(a) Performance on Random instances
 Scenario S-R1A1 S-R5A1 S-R5A3 D-R2A2O2 D-R5A2O2 D-R5A3O1
@@
 - 2.55 - 1.43 - 0.90 - 2.82 - 1.33 - 0.79
-(b) Performance on 100% set with and without repacking.
-We define six distinct scenarios to evaluate our bin packing framework under varying robotic configurations and perception constraints, as summarized in Table I and shown in Fig.
+(b) Performance on 100% set with and without repacking
+V-C Test Environment
+V-C1 Scenarios
+We define six distinct scenarios to evaluate the proposed bin packing framework under varying robotic configurations and perception constraints, as summarized in Table II and illustrated in Fig.
 1.
-Each scenario label encodes the following:
+Each scenario label is encoded as follows:
   • S (single robot) or D (dual robot): robot configuration,
   • R: maximum number of recognized items,
-  • A: number of accessible items per robot,
-  • O: number of items in the overlapping workspace (for dual-robot settings only).
-Scenarios S-R1A1, S-R5A1, and S-R5A3 represent single-arm configurations, while D-R2A2O2, D-R5A2O2, and D-R5A3O1 correspond to dual-arm settings.
-S-R1A1 denotes a minimal setup where one robot can recognize and access only a single item.
-In contrast, D-R5A3O1 reflects the most capable configuration, where two robots can recognize five items, each access three, and share one item in the overlapping region.
-V-C2 Test with Random Instances
+  • A: maximum number of accessible items per robot,
+  • O: maximum number of items in the overlapping workspace (for dual-robot settings only).
+Scenarios S-R1A1, S-R5A1, and S-R5A3 represent single-robot configurations, while D-R2A2O2, D-R5A2O2, and D-R5A3O1 correspond to dual-robot settings.
+S-R1A1 denotes the minimal case, where a single robot can recognize and access only one item at a time.
+In contrast, D-R5A3O1 represents the most capable configuration, where two robots can recognize five items, each can access three items, and one item lies within the overlapping workspace shared by both robots.
+The maximum number of items that can be accessed by all robots, denoted as $n_{\max}$, is given by $n_{\max}=n_{\text{A}}\times n_{\text{robot}}-n_{\text{O}}\times(n_{\text{robot}}-1)$, where $n_{\text{A}}$ denotes the number of accessible items per robot, $n_{\text{O}}$ represents the number of items located in the overlapping workspace, and $n_{\max}\leq n_{\text{R}}$, with $n_{\text{R}}$ being the number of recognized items.
+V-C2 Conveyor Logic
+Since our search algorithm considers both robot accessibility and task allocation, the conveyor belt logic for positioning incoming items is a critical component of the experimental setup.
+For single-robot scenarios or dual-robot configurations where both robots share the same accessible region (e.g., $n_{\text{A}}=n_{\text{O}}$ in D-R2A2O2 and D-R5A2O2), items can be continuously positioned along the conveyor.
+However, in configurations such as D-R5A3O1, where the maximum number of items accessible to each robot ($n_{\text{A}}$) differs from that in the overlapping region ($n_{\text{O}}$), exclusive zones arise that can be reached by only one robot.
+If more items accumulate in one exclusive zone, the advantages of using dual robots diminish due to workload imbalance.
+To maintain balanced item allocation, we model the conveyor control logic at each time step as follows.
+Notation:
+  • $k_{\text{total}}$: Total number of items on the conveyor
+  • $k_{\alpha}$: Number of items in the exclusive region of the rear robot
+  • $k_{\beta}$: Number of items in the exclusive region of the front robot
+  • $k_{\text{O}}$: Number of items in the overlapping region (accessible to both robots)
+  • $t$: Time step before conveyor operation
+  • $t+1$: Time step after conveyor operation
+When $k_{\text{total},t+1}\geq n_{\text{max}}$, where $n_{\text{max}}=2n_{\text{A}}-n_{\text{O}}$ denotes the maximum number of items that both robots can collectively access, items are continuously positioned along the conveyor as in the fully shared-region cases.
+When $k_{\text{total},t+1}<n_{\text{max}}$, the item distribution across regions is determined as follows:
+$k_{\text{O},t+1}$ $=\min\!\left(k_{\text{total},{t+1}}-k_{\alpha,t},\,n_{\text{O}}\right),$ (7)
+$k_{\alpha,{t+1}}$ $=\max\!\left(k_{\alpha,t},\,\left\lfloor\frac{k_{\text{total},t+1}}{2}\right\rfloor+1-k_{\text{O},t+1}\right),$ (8)
+$k_{\beta,{t+1}}$ $=k_{\text{total},t+1}-k_{\alpha,t+1}-k_{\text{O},t+1}.$ (9)
+By controlling the conveyor belt to place items sequentially into the $\alpha$, O, and $\beta$ regions according to the computed quantities, the system maintains a balanced allocation of items between the workspaces of the two robots even when their accessible regions differ, thereby preserving the efficiency gains of dual-robot operation.
+All following tests are conducted under the assumption that the conveyor belt operates according to the described logic at every time step.
+V-D Algorithm Tests
+Before discussing each result in detail, we briefly summarize the purpose and scope of the ablation studies conducted to verify the contribution of each component in the proposed hierarchical framework.
+In summary, the ablation results demonstrate that:
+  • under the consideration of rotation and multiple perceived items, the tree search component improves orientation selection and packing efficiency;
+  • the dual-manipulator setting significantly reduces packing steps while maintaining high utilization;
+  • the tree search handles multiple reachable items and performs task allocation, especially in dual-manipulator settings;
+  • the repacking algorithm further enhances bin utilization across all scenarios within a limited planning time;
+  • the reinforcement learning module alone achieves higher utilization than the baseline algorithms.
+Detailed analyses of each case are presented in the following subsections.
+V-D1 Test with Random Instances
 We evaluate our method on the Random instances with 1000 episodes per scenario, where each episode involves packing a single bin.
+In the Random instance, the items that has not been packed at the current episode, are not discarded but are reconsidered in subsequent planning stages.
+Such items are revisited either during the repacking phase or when planning for the next bin, ensuring that all items have the opportunity to be packed.
+This cumulative process over 1,000 bins is reflected in the total number of packed items reported in Table III(a).
 All scenarios share the same item arrival sequence across the entire test.
 We first evaluate our method without repacking across six scenarios, with and without item rotation.
-As shown in Table II(a), both bin utilization and the total number of packed items consistently improve when rotation is enabled.
-For instance, in Scenario S-R1A1, the utilization increases from 78.88% to 85.17%, and the number of packed items from 8715 to 9422.
+When performing packing only, the Selection described in Alg.
+2 (line 18) is skipped, and the total planning time for all scenarios averages less than 1.2 seconds.
+As shown in Table III(a), both bin utilization and the total number of packed items consistently improve when rotation is enabled.
+For instance, in Scenario S-R1A1, the utilization increases from 78.88% to 85.17%, and the number of packed items from 8715 to 9422, achieving higher throughput within the same number of bins.
 Similar trends are observed across all scenarios, with Scenario D-R5A3O1 achieving the highest utilization of 98.59%.
+Furthermore, the influence of the number of perceived items on performance is verified by comparing S-R1A1 vs.
+S-R5A1 and D-R2A2O2 vs.
+D-R5A2O2.
+In addition, when comparing S-R5A1 vs.
+S-R5A3 and D-R5A2O2 vs.
+D-R5A3O1, where each robot has more reachable objects, the results show consistent improvements in utilization.
+In the dual-arm cases (D-R5A2O2 vs.
+D-R5A3O1), even with a larger number of packed items, the total number of packing steps decreased.
+These results collectively verify that the proposed candidate-selection strategy is valid and that the method successfully performs task allocation between the two manipulators, directly indicating improved throughput per cell.
 We then test repacking in four scenarios where there is room for improvement: S-R1A1, S-R5A1, D-R2A2O2, and D-R5A2O2.
+In these repacking experiments, the number of candidates selected by the Selection function is fixed to two to compare the utilization achieved under a limited additional planning time for repacking.
 Fig.
@@
 In Scenario S-R5A1, the utilization improves from 92.38% to 99.20%, and in D-R5A2O2 from 96.46% to 99.67%.
+The fact that such high improvements are achieved within a short planning time indicates that the proposed last-packed-first search strategy efficiently identifies promising repacking policies.
 Scenarios with more known items result in fewer repacking actions, which helps reduce motion overhead during execution.
@@
 (b) The total number of repacked items increases with longer planning time.
-V-C3 Test with 100% set
-We evaluate our method on the 100% set with 3000 episodes per scenario, where each episode uses the same fixed item arrival sequence.
-As shown in Table II(b), our method achieves high utilization even without repacking.
-When repacking is enabled, 100% utilization is achieved in all scenarios.
+Figure 9: Repacking success count and planning time with varying numbers of candidates selected by the Selection in D-R5A3O1 on the 100% set.
+V-D2 Test with 100% set
+We evaluate our method on the 100% set with 3000 episodes, where each episode uses the same fixed item arrival sequence per scenario.
+In the 100% set experiments, items that have no valid placement are re-evaluated during the repacking phase, and if infeasible placements remain, those items are excluded from further consideration, since the item sets are predefined for each bin.
+When performing packing only, the Selection in Alg.
+2 (line 18) is skipped, and the total planning time for all scenarios averages less than 1.3 seconds.
+As shown in Table III(b), our method achieves high utilization even without repacking.
+Furthermore, performance improves as the number of perceived and reachable items increases, confirming that the hierarchical decision process leverages additional information.
+In addition, we conduct an experiment on the 100% set to measure the repacking success count and computation time with varying numbers of candidates selected by the Selection function in D-R5A3O1, as presented in Fig.
+9.
+The Selection process is applied until the number of remaining items in the search tree drops below three; when fewer than three items remains, pruning is skipped to preserve remaining candidates, since the search space becomes sufficiently small due to additional early-termination conditions implemented in Alg.2 (lines 10–11,29–30).
+When repacking is enabled with this setting, 100% utilization is achieved across all scenarios.
 The number of fully packed bins increases significantly with repacking, from 684 to 3000 in S-R1A1 and from 865 to 3000 in D-R2A2O2.
 Repacking improves packing quality at the cost of moderate computation time, with an average of less than 3 seconds to find a repacking solution.
-V-C4 Compare with Baseline
-TABLE III: Comparison of bin utilization (%) with baseline methods in $4\times 4$ and $5\times 5$ bin environments.
+These findings further confirm that the proposed hierarchical framework achieves near-optimal utilization under realistic time constraints.
+V-D3 Compare with Baseline
+TABLE IV: Comparison of bin utilization (%) with baseline methods in $4\times 4$ and $5\times 5$ bin environments.
 Method Uti. (%)
@@
 The evaluation follows the same setup as the baselines in $4\times 4$ and $5\times 5$ bin environments under a fully online setting without rotation or repacking, ensuring fair comparison.
-Table III shows that our method achieves the highest mean bin utilization in both settings—98.23% in the $4\times 4$ bin (compared to 94.87% for Deep-Pack and 97.23% for BERM), and 96.58% in the $5\times 5$ bin (compared to 91.48% and 94.30%, respectively).
-These results demonstrate that our DRL-based policy outperforms the state-of-the-art methods in a constrained online scenario.
-V-D Simulation Tests
+Table IV shows that our method achieves the highest mean bin utilization in both settings—98.23% in the $4\times 4$ bin (compared to 94.87% for Deep-Pack and 97.23% for BERM), and 96.58% in the $5\times 5$ bin (compared to 91.48% and 94.30%, respectively).
+Assuming that each pallet or tote corresponds to a fixed container volume, this improvement directly translates to a reduction of roughly one pallet for every 20–40 bins packed, resulting in measurable logistics cost savings and reduced storage requirements.
+The superior performance of our method can be attributed to the DRL formulation.
+Compared with heuristic-based baselines [5, 4], which usually follow sequential placement heuristics, our RL agent is free to explore and select placements from any direction within the bin, allowing more flexible decision-making.
+Furthermore, our method surpasses existing DRL-based algorithms [7, 26] due to two main design choices:
+  • the use of diverse training datasets that improve generalization across different item distributions,
+  • a reward formulation defined as the number of occupied cells adjacent to the packed item (described in Sec.
+IV-A), which provides dense and flexible rewards and induces generalizable packing behaviors.
+V-E Simulation Tests
+V-E1 Configuration of Simulation Environment
 The simulation environment is developed using Unity [30] to evaluate the proposed framework under physical constraints.
 Universal Robots UR5e is used as the manipulator in both single- and dual-arm configurations.
-The simulation environments for each setting are shown in Fig.
-9.
-For motion planning, cuRobo [29] is employed in both configurations, while StopNGo [13] is integrated to coordinate dual-arm execution.
+A square bin of 45 cm $\times$ 45 cm is used, discretized into a $10~\times~10$ grid where each cell corresponds to 4.5 cm.
+The simulation environment layout for the dual-robot scenario, and the packing scenes of single-arm and dual-arm configurations are illustrated in Fig.
+11.
+The design of the simulation environment is primarily guided by the workspaces of both manipulators.
+Each robot is positioned such that both arms can fully cover the bin area while maintaining sufficient reach to grasp incoming items on the conveyor belt.
+To prevent mutual motion interference, the robots are spaced approximately half of their reach distance apart, which reflects safety and accessibility considerations.
+For the single-robot configuration, only the right arm is used to ensure consistent comparison under identical environmental conditions.
+Fig.
+10 illustrates the action sequence of the robot during the packing process.
+The interval between incoming items on the conveyor belt is set to 0.7 m, and the conveyor speed is configured to 0.5 m/s.
+When the conveyor operates and a new item arrives, the planning module generates a placement plan, after which the robot initiates the packing task.
+The packing operation consists of repeated pick-and-place actions.
+Once the picking of the current item on the conveyor is completed, the conveyor operates again to transport the next item, thereby allowing conveyor operation and robotic placement to proceed in parallel.
+Once the new item has arrived and the planning for the next step is completed, followed by the completion of the previous placement, the next operation begins (if repacking is required, the repacking phase is executed first).
+V-E2 Makespan Evaluation
+In this paper, we define the makespan as the total time required for the robot(s) to complete the loading process within a single episode, including conveyor operation, planning, and execution, as shown in Fig.
+10.
+For motion planning, cuRobo [31] is employed in both configurations, while Stop-N-Go [13] is integrated to coordinate dual-arm execution.
 Each setup is tested over 30 episodes using the 100% set instances.
-As shown in Table IV, the dual-arm scenario achieves a shorter makespan than the single-arm scenario, despite requiring longer per-step planning durations.
-Specifically, it reduces the overall makespan by approximately 24% in both repacking and non-repacking cases.
-Since planning and execution proceed concurrently, planning for the next step is typically completed during the current execution, minimizing idle time.
+The average time between the conveyor activation and the arrival of a new item is approximately 1.7 seconds.
+As shown in Table V, the dual-arm scenario achieves a shorter makespan than the single-arm scenario, despite requiring longer per-step planning durations.
+Specifically, it reduces the overall makespan by approximately 24% in both repacking and non-repacking cases, corresponding to a proportional increase in cell throughput under identical task conditions.
+Since planning and execution proceed concurrently, planning for the next step is typically completed during the current execution, thereby minimizing idle time.
 Although repacking introduces slight overhead in planning and execution, it improves bin utilization (as discussed in Sec.
-V-C3), which justifies the additional cost.
+V-D2), which justifies the additional cost.
 In instances with an odd number of items, the final item is always handled by a single robot, limiting the relative gain in makespan reduction.
 These results demonstrate that the proposed framework can generate physically feasible and time-efficient plans for dual-arm robotic systems.
-(a) S-R5A3 simulation environment
-(b) D-R5A3O1 simulation environment
-Figure 9: Simulation environments used for testing. (a) single-arm setting (S-R5A3), and (b) dual-arm setting (D-R5A3O1).
-TABLE IV: Planning and Execution Time in Single- and Dual-Arms
+V-F Physical Robot Test
+In the physical robot experiments, a 32 cm $\times$ 26 cm bin is used with 1 cm margins on both sides, corresponding to a 3 cm grid resolution.
+Two AgileX Piper manipulators [32] are employed for dual-arm operation, following the D-R5A3O1 scenario setting.
+ArUco markers are attached to the objects for vision-based detection and pose estimation.
+The experimental setup and dual-arm operation are illustrated in Fig.
+12.
+Motion planning for dual manipulators is performed using cuRobo [31] and Stop-N-Go [13].
+Through these experiments, we demonstrated the feasibility of applying the proposed framework to real dual-arm packing tasks in a logistics environment.
+Figure 10: Action sequence and makespan of single-arm and dual-arm configurations.
+(a) Layout of the dual-arm simulation environment (D-R5A3O1)
+(b) Single-arm packing scene (S-R5A3)
+(c) Dual-arm packing scene (D-R5A3O1)
+Figure 11: Overview of the simulation environment and packing scenes.
+In (a), the layout of the dual-arm simulation environment is shown, where the circles indicate the workspace of each manipulator, while (b) and (c) show the packing scenes of the single-arm and dual-arm configurations, respectively.
+TABLE V: Planning and Execution Time in Single- and Dual-Arms.
 Scenario S-R5A3 D-R5A3O1
@@
 124.48 130.98 94.29 100.12
+Figure 12: Real dual-arm packing experiment setup showing the manipulators, bin, and vision-based detection system.
 VI Conclusion
@@
 These results highlight the benefits of repacking and parallel execution for practical robotic systems.
+Future work is left for extending the current 2D framework to 3D by replacing the binary grid matrix representation with a height-map formulation, while incorporating dual-arm workspace constraints for three-dimensional packing tasks.